Two elements X and Y (atomic mass of x is equal to 75 and Y is equal to 16) combined to give a compound having 76% of x. The formula of compound is

Grade:12

1 Answers

ANIKET RAJ
5 years ago
Easy one..................
apply formula for avg.molar mass= A*a%+B*b%/100
let %of X be a%
tne % of Y=100-a
75*a+16*(100-a)/100=76%of 75
59a+1600=76*75
a=70 100-a =30
ratio =70/75=0.9
ratio=30/16=1.8
formula=XY2